Write your first derivative of the following function Y = 2X3 -3X2 +2X -50
Answer #1

\frac{dy}{dx}=\frac{d}{dx}(2x^3-3x^2+2x-50)

=6x2 -6x +2

\frac{dy}{dx}kx^n=(n*k)x^{n-1}